# Sequential Rank and Preference Learning with the Bayesian Mallows Model

### v1

@@ Abstract
The Bayesian Mallows model is a flexible tool for analyzing data in the form of complete or partial rankings, and transitive or intransitive pairwise preferences. In many potential applications of preference learning, data arrive sequentially and it is of practical interest to update posterior beliefs and predictions efficiently, based on the currently available data. Despite this, most algorithms proposed so far have focused on batch inference. In this paper we present an algorithm for sequentially estimating the posterior distributions of the Bayesian Mallows model using nested sequential Monte Carlo. As it requires minimum user input in form of tuning parameters, is straightforward to parallelize, and returns the marginal likelihood as a direct byproduct of estimation, the algorithm is an alternative to Markov chain Monte Carlo techniques also in batch estimation settings.
Keywords: Mallows mixtures, partial rankings, particle filter, preference learning, SMC2.
@@ 1 Introduction
The paper proceeds as follows. In Section 2 we provide necessary background on the Bayesian Mallows model. In Section 3 we propose an SMC2 algorithm for the Bayesian Mallows model with partial rankings or pairwise preference data. In Section 4 we report the results of simulation experiments demonstrating the algorithm. We conclude and discuss further developments in Section 5. An R (R Core Team,, 2024) package providing an API to our C++ implementation is available from GitHub [^1] https://github.com/osorensen/BayesMallowsSMC2 and R code for reproducing all the results in the paper is available from our OSF repository. [^2] https://osf.io/pquk4/
@@ 2.3 Distance Functions and Normalizing Constants
Computing the normalizing constant $Z(\alpha)$ as in (2) requires summing over $|\mathcal{P}_{m}|=m!$ terms, but tractable exact expressions exist for Cayley, Kendall, and Hamming distances (Fligner and Verducci,, 1986; Irurozki et al.,, 2018). Furthermore, since $d(\bm{r},\bm{e})$ takes on a set of $l<m!$ values $\mathcal{D}=\{d_{1},d_{2},\dots,d_{l}\}$, we can define $L_{i}=\{\bm{r}\in\mathcal{P}_{m}:d(\bm{r},\bm{e})=d_{i}\}$ and write $Z(\alpha)=\sum_{i=1}^{l}|L_{i}|e^{-\alpha d_{i}}$. For the footrule, $l=\mathcal{O}(m^{2})$, for Spearman distance $l=\mathcal{O}(m^{3})$, and for Ulam distance $l=\mathcal{O}(m)$ (Crispino,, 2018; Crispino et al.,, 2023; Irurozki,, 2014; Irurozki et al.,, 2016). Unfortunately, while the set of distances $\mathcal{D}$ is well known, finding the cardinalities $|L_{i}|$ is hard. The Online Encyclopedia of Integer Sequences (Sloane,, 2023) contains $|L_{i}|$ up to $m=50$ for the footrule, up to $m=20$ for Spearman distance, and $m=60$ for Ulam distance. Beyond these upper limits, asymptotic approximations exist for the footrule and Spearman distances (Crispino et al.,, 2023; Mukherjee,, 2016), and an importance sampling scheme has been developed by Vitelli et al., (2017). The latter can in principle be run to arbitrary precision, and importantly, estimates of the partition function can be precomputed for a given number of items $m$ over a grid of $\alpha$ values (Sørensen et al.,, 2020). Thus, in the rest of this paper we assume that $Z(\alpha)$ is available and that its Monte Carlo error (if any) is negligible compared to the Monte Carlo error of the proposed SMC2 algorithm.
@@ 3.4.2 Consistent Pairwise Preferences
We will sample latent rankings uniformly on $\text{TO}_{n}=\text{TO}(\bm{y}_{n})$, the set of topological orderings of $\bm{y}_{n}$, and hence need to both count and generate linear extensions. The counting problem itself is known to be #P complete (Brightwell and Winkler,, 1991), although faster algorithms exist for special cases, e.g., sparse posets (Kangas et al.,, 2016). Generation of the linear extensions can be obtained in constant additional time (Pruesse and Ruskey,, 1994), and very compact storage of the extensions can be obtained using Gray codes (Ono and Nakano,, 2005; Pruesse and Ruskey,, 1994) or permutation decision diagrams (Inoue and Minato,, 2014). In our implementation we used depth-first search (Cormen et al.,, 2022, Ch. 20.3-20.4), and generated all orderings by looping over all child nodes at each recursive step of the algorithm, keeping track of the solutions via backtracking. To keep the implementation relatively simple, we stored each element of $\text{TO}_{n}$ to disk and randomly sampled among the files when proposing a new latent ranking.
@@
4 Simulation Experiments
We here report results of simulation experiments aimed at testing the proposed algorithms. Simulations were performed on a MacBook Pro with a 32 GB Apple M1 Max chip, and on the high performance computing cluster Fox, provided by the University of Oslo. The R packages Rcpp (Eddelbuettel and François,, 2013) and RcppArmadillo (Eddelbuettel and Sanderson,, 2014) were used as interfaces to the C++ implementation of our algorithms, which made heavy use of the Armadillo library (Sanderson and Curtin,, 2016). Our parallelization of SMC2 used the futures framework (Bengtsson,, 2021), through the furrr package (Vaughan and Dancho,, 2021). Pre- and post-processing of data, as well as visualization, was done mainly with the set of R packages provided by the tidyverse (Wickham et al.,, 2019).
4.1 Generation of Topological Orderings
The generation of all topological orderings when proposing latent rankings in the pairwise preference case is a potential bottleneck. We here report two numerical experiments investigating the extent of this issue in real data. In both experiments we defined $\text{TO}_{n}$ as the number of orderings of the items compared by user $n$, since permuting the non-compared items is a computationally easy task.
4.1.1 Topological Orderings for PrefLib Data
@@
4.1.2 Topological Orderings for Beach Preference Data
@@ 4.2 Complete Rankings
To study the performance of SMC2 in a sequential inference case with complete rankings, we generated 100 datasets with complete rankings from the Mallows model with the footrule distance using the sampling algorithm of Vitelli et al., (2017, Appendix C). In all simulations there were $m=10$ items and $N=1000$ users, the scale parameter was $\alpha=0.1$ and there was a single cluster. The users were assumed to enter one at a time, yielding 1,000 timepoints. For all simulated datasets the algorithm was run in parallel on $P$ cores, with $P\in\{1,2,4,6,8\}$ and $R=5000/P$ particles, using multinomial, residual, stratified, and systematic resampling. The gamma prior for $\alpha$ had shape $\gamma=1$ and rate $\lambda=0.5$ and the resampling threshold was $A=R/2$. Since the likelihood increments are analytically given by (14), the number of particle filters was fixed to 1 and the doubling threshold $B$ in Algorithm 3 set to 00.
The simulation results are summarized in Figure 3. The left plot illustrates how the computing time depends on the resampling scheme and parallelization. Regarding the former, all four resampling schemes performed equally fast. Furthermore, the plot shows a clear benefit of parallelization, although with slightly diminishing returns: doubling the number of cores from 1 to 2, 2 to 4, and 4 to 8, respectively, led to factors 1.85, 1.81, and 1.67 speed-up, respectively. The average time with eight cores was 25.9 seconds. For comparison, batch estimation with MCMC with a burnin-in period of 1,000 iterations and 5,000 post-burnin iterations took on average 3 seconds. However, sequential estimation with MCMC was prohibitively slow, as the algorithm has to be rerun at each timepoint; just the last ten timepoints from 991 to 1000 took more time than the complete sequential inference with SMC on eight cores.
@@ 4.3 Top-$k$ Rankings
We next considered the case in which users provide top-3 rankings of $m=10$ items. As in Section 4.2, the users were assumed to arrive sequentially, one at each timepoint. A total of 30 random datasets were simulated, with $N=200$ users, scale parameter $\alpha=0.3$, and footrule distance. Since each user had only ranked three out of ten items, the particle filter in Algorithm 1 now had to be run to integrate over the remaining seven items for each user. The total number of particles was set to $R=3000$ which were processed separately on eight cores and combined using (17). The resampling threshold was $A=R/2$ and the threshold for particle filter doubling was set to an average acceptance rate $B=0.2$ in the rejuvenation step. The initial number of particle filters per core was set to $S=20$. Both uniform and pseudolikelihood proposals were used.
@@
Table 1: Results of simulations with sequentially arriving top-3 rankings. Values are Monte Carlo averages, with standard errors in parentheses.
Proposal Resampler Posterior mean $\alpha$ Posterior mean $d(\bm{\rho},\hat{\bm{\rho}})$
Pseudolikelihood Multinomial 0.299 (9e-04) 3.09 (0.058)
Pseudolikelihood Residual 0.294 (9e-04) 2.95 (0.055)
Pseudolikelihood Stratified 0.301 (9e-04) 2.96 (0.062)
Pseudolikelihood Systematic 0.296 (9e-04) 2.98 (0.054)
Uniform Multinomial 0.298 (7e-04) 3.19 (0.047)
Uniform Residual 0.296 (8e-04) 3.55 (0.049)
Uniform Stratified 0.293 (9e-04) 3.46 (0.048)
Uniform Systematic 0.295 (7e-04) 3.07 (0.048)
@@
Table 1 shows simulation results after the final timepoint. The third column shows that the final estimate of $\alpha$ was close to the true value for all algorithm settings. With uniform latent rank proposal, systematic resampling gave slightly lower posterior mean distance to the true modal ranking $\bm{\rho}$ than the other alternatives. Overall, pseudolikelihood proposal gave posterior means of $\bm{\rho}$ closer to the true ranking. Processing a single dataset on eight cores took about 12 minutes when using uniform proposal, and between 10 and 100 times more with pseudolikelihood proposal. [^5] Due to the longer computing time, the models with pseudolikelihood proposal were run on the Fox computing cluster, while the models with uniform proposal were run on the MacBook Pro. Thus, the computing times are not exactly comparable. Hence, the increase in accuracy of $\bm{\rho}$ has a cost in terms of computational demand.
@@
Figure 4: Number of particle filters and resampling probability for simulations with sequentially arriving top-3 rankings. Shaded regions are 95% confidence bands.
@@
Figure 4 shows the number of particle filters as a function of time (left) and the resampling probability as a function of time (right), with uniform proposal distribution. The plots for pseudolikelihood proposal were almost identical, and are not shown. The curves were obtained by fitting generalized additive models (GAMs) (Wood,, 2017) with ten thin-plate regression splines (Wood,, 2003) as basis functions to the number of particle filters and a binary resampling indicator, respectively. A unit link function was used for the number of particles and a logit link for the resampling indicator, and smoothing was done with restricted maximum likelihood. The left plot shows that the number of particles grows close to linearly. This is expected from Theorem 1 of Andrieu et al., (2010), which implies that the number of particle filters must grow linearly with the number of observations for the acceptance rate to stay constant. The right plot shows that the resampling probability decreases with the number of timepoints. Since each resampling step is followed by one or more rejuvenation steps, this means that as rejuvenation becomes more computationally demanding due to more observations, it also becomes less frequent. This behavior also agrees with what is expected from theory; in particular, Theorem 1 and Section 4.3 of Chopin, (2002) predicts that the time interval between each time when resampling is needed should increase geometrically in the total number of observations, which would produce an exponentially decaying curve (see also Proposition 17.1 in Chopin and Papaspiliopoulos,, 2020).
@@ 4.4 Pairwise Preferences
We next simulated consistent pairwise preference data by first generating complete rankings of $m=5$ items with $\bm{\rho}=(1,2,\dots,5)^{\prime}$ and $\alpha=0.3$ for 200 users, and then randomly selecting four implied pairwise preferences for each user. The users were assumed to arrive sequentially, one at each timepoint. The process was repeated 30 times, and each dataset was processed in parallel on 10 CPUs, each with $S=300$ particles. The initial number of particle filters was 20, and all other parameters were as described in the previous sections.
Across the 30 datasets, the average posterior mean of $\alpha$ at the final timepoint was $0.325$, with 95% Monte Carlo interval $(0.314,0.366)$. For comparison, the average posterior mean of $\alpha$ obtained by MCMC estimation with 20,000 post-burnin iterations on the same datasets was $0.300$ with 95% Monte Carlo interval $(0.289,0.310)$, suggesting that SMC2 was slightly positively biased in this case. Figure 5 (left) shows how the posterior expectation of $\alpha$ evolved as more data became available, with initial rapid fluctuations followed by a stabilization close to the true value of $0.3$ after about 100 timepoints. The center plot in Figure 5 shows the posterior probability that item 1 is preferred to item 2 in the modal ranking, and the right plot show the posterior probability that item 1 is preferred to item 3. As expected, the latter converges more quickly to one than the former. In particular, for three of the simulated datasets $P(\rho_{1}<\rho_{2})$ was below 0.5 even after the final timepoint, implying that items 1 and 2 were "flipped" in the posterior distribution, while $P(\rho_{1}<\rho_{3})$ was close to 1 for all datasets.

### v2

@@ Abstract
The Bayesian Mallows model is a flexible tool for analyzing data in the form of complete or partial rankings, and transitive or intransitive pairwise preferences. In many potential applications of preference learning, data arrive sequentially and it is of practical interest to update posterior beliefs and predictions efficiently, based on the currently available data. Despite this, most algorithms proposed so far have focused on batch inference. In this paper we present an algorithm for sequentially estimating the posterior distributions of the Bayesian Mallows model using nested sequential Monte Carlo. The algorithm requires minimal user input in the form of tuning parameters, is straightforward to parallelize, and returns the marginal likelihood as a direct byproduct of estimation. We evaluate its performance in simulation experiments, and illustrate a real use case with sequential ranking of Formula 1 drivers throughout three seasons of races.
@@ 1 Introduction
A motivating example is shown in Figure 1, showing race results from the Formula 1 seasons 2022-2024.111Data were downloaded from https://github.com/toUpperCase78/formula1-datasets. We included the 16 drivers who completed 50 or more of the total 68 races, and computed rankings from the results of each race. If we assume that the underlying skills of the drivers – as well as the performance of the cars and teams – are relatively stable across the three seasons, the results of each race can be viewed as a noisy assessment of the underlying true ranking. Using the SMC2 algorithm proposed in this paper we can easily update the posterior distribution for the ranking of the drivers after each race. We will revisit these data in Section 6.
Figure 1: Rankings of drivers based on race results in Formula 1 seasons 2022, 2023, and 2024. Vertical dashed lines indicate the start of a season. Traces for the three drivers with best mean ranking across all races are shown in color.
The paper proceeds as follows. In Section 2 we provide necessary background on the Bayesian Mallows model. In Section 3 we propose an SMC2 algorithm for the Bayesian Mallows model with partial rankings or pairwise preference data. In Section 4 we investigate the computational requirements for computing topological orderings – an essential part of the algorithms. In Section 5 we report the results of simulation experiments and in Section 6 we revisit the Formula 1 data. We conclude and discuss further developments in Section 7. An R (R Core Team,, 2024) package providing an API to our C++ implementation is available from GitHub [^2] https://github.com/osorensen/BayesMallowsSMC2 and R code for reproducing all the results in the paper is available from our OSF repository. [^3] https://osf.io/pquk4/
@@ 2.3 Distance Functions and Normalizing Constants
Computing the normalizing constant $Z(\alpha)$ as in (2) requires summing over $|\mathcal{P}_{m}|=m!$ terms, but tractable exact expressions exist for Cayley, Kendall, and Hamming distances (Fligner and Verducci,, 1986; Irurozki et al.,, 2018). Furthermore, since $d(\bm{r},\bm{e})$ takes on a set of $l<m!$ values $\mathcal{D}=\{d_{1},d_{2},\dots,d_{l}\}$, we can define $L_{i}=\{\bm{r}\in\mathcal{P}_{m}:d(\bm{r},\bm{e})=d_{i}\}$ and write $Z(\alpha)=\sum_{i=1}^{l}|L_{i}|e^{-\alpha d_{i}}$. For the footrule, $l=\mathcal{O}(m^{2})$, for Spearman distance $l=\mathcal{O}(m^{3})$, and for Ulam distance $l=\mathcal{O}(m)$ (Crispino,, 2018; Crispino et al.,, 2023; Irurozki,, 2014; Irurozki et al.,, 2016). Unfortunately, while the set of distances $\mathcal{D}$ is well known, finding the cardinalities $|L_{i}|$ is hard. The Online Encyclopedia of Integer Sequences (Sloane,, 2023) contains $|L_{i}|$ up to $m=50$ for the footrule, up to $m=20$ for Spearman distance, and $m=60$ for Ulam distance. Beyond these upper limits, asymptotic approximations exist for the footrule and Spearman distances (Crispino et al.,, 2023; Mukherjee,, 2016), and an importance sampling scheme has been developed by Vitelli et al., (2017). The latter can in principle be run to arbitrary precision, and importantly, estimates of the partition function can be precomputed for a given number of items $m$ over a grid of $\alpha$ values (Sørensen et al.,, 2020). Thus, in the rest of this paper we assume that $Z(\alpha)$ is available and that its Monte Carlo error (if any) is negligible compared to the Monte Carlo error of the proposed SMC2 algorithm. In the simulations and application examples, the number of items $m$ is always such that $Z(\alpha)$ is known exactly.
@@ 3.4.2 Consistent Pairwise Preferences
We will sample latent rankings uniformly on $\text{TO}_{n}=\text{TO}(\bm{y}_{n})$, the set of topological orderings of $\bm{y}_{n}$, and hence need to both count and generate linear extensions. The counting problem itself is known to be #P complete (Brightwell and Winkler,, 1991), although faster algorithms exist for special cases, e.g., sparse posets (Kangas et al.,, 2016). Generation of the linear extensions can be obtained in constant additional time (Pruesse and Ruskey,, 1994), and very compact storage of the extensions can be obtained using Gray codes (Ono and Nakano,, 2005; Pruesse and Ruskey,, 1994) or permutation decision diagrams (Inoue and Minato,, 2014). In our implementation we used depth-first search (Cormen et al.,, 2022, Ch. 20.3-20.4), and generated all orderings by looping over all child nodes at each recursive step of the algorithm, keeping track of the solutions via backtracking.
@@
4 Generation of Topological Orderings
The generation of all topological orderings when proposing latent rankings in the pairwise preference case is a potential bottleneck. We here report two numerical experiments investigating the extent of this issue in real data. In both experiments we defined $\text{TO}_{n}$ as the number of orderings of the items compared by user $n$, since permuting the non-compared items is a computationally easy task. Computations were performed on a MacBook Pro with a 32 GB Apple M1 Max chip. Further details about the implementation are provided in the first paragraph of Section 5.
4.1 Topological Orderings for PrefLib Data
@@
4.2 Topological Orderings for Beach Preference Data
@@
5 Simulation Experiments
We here report results of simulation experiments aimed at testing the proposed algorithms. The R packages Rcpp (Eddelbuettel and François,, 2013) and RcppArmadillo (Eddelbuettel and Sanderson,, 2014) were used as interfaces to the C++ implementation of our algorithms, which made heavy use of the Armadillo library (Sanderson and Curtin,, 2016). Our parallelization of SMC2 used the futures framework (Bengtsson,, 2021), through the furrr package (Vaughan and Dancho,, 2021). Pre- and post-processing of data, as well as visualization, was done mainly with the set of R packages provided by the tidyverse (Wickham et al.,, 2019).
@@ 5.1 Complete Rankings
To study the performance of SMC2 in a sequential inference case with complete rankings, we generated 100 datasets with complete rankings from the Mallows model with the footrule distance using the sampling algorithm of Vitelli et al., (2017, Appendix C). In all simulations there were $m=10$ items and $N=1000$ users, the scale parameter was $\alpha=0.1$ and there was a single cluster. The users were assumed to enter one at a time, yielding 1,000 timepoints. For all simulated datasets the algorithm was run in parallel on $P$ cores, with $P\in\{1,2,4,6,8\}$ and $R=5000/P$ particles, using multinomial, residual, stratified, and systematic resampling. The gamma prior for $\alpha$ had shape $\gamma=1$ and rate $\lambda=0.5$ and the resampling threshold was $A=R/2$. Since the likelihood increments are analytically given by (14), the number of particle filters was fixed to 1 and the doubling threshold $B$ in Algorithm 3 set to 00. Simulations were run on the MacBook Pro mentioned in Section 4.
The simulation results are summarized in Figure 4. The left plot illustrates how the computing time depends on the resampling scheme and parallelization. Regarding the former, all four resampling schemes performed equally fast. Furthermore, the plot shows a clear benefit of parallelization, although with slightly diminishing returns: doubling the number of cores from 1 to 2, 2 to 4, and 4 to 8, respectively, led to factors 1.93, 1.82, and 1.76 speed-up, respectively. The average time with eight cores was 16.5 seconds. For comparison, batch estimation with MCMC with a burnin-in period of 1,000 iterations and 5,000 post-burnin iterations took on average 3.2 seconds.
@@ 5.2 Top-$k$ Rankings
We next considered the case in which users provide top-3 rankings of $m=10$ items. As in Section 5.1, the users were assumed to arrive sequentially, one at each timepoint. A total of 80 random datasets were simulated, with $N=200$ users, scale parameter $\alpha=0.3$, and footrule distance. Since each user had only ranked three out of ten items, the particle filter in Algorithm 1 now had to be run to integrate over the remaining seven items for each user. The total number of particles was set to $R=3000$ which were processed in parallel on 20 cores and combined using (17). The resampling threshold was $A=R/2$ and the threshold for particle filter doubling was set to an average acceptance rate $B=0.2$ in the rejuvenation step. The initial number of particle filters per core was set to $S=20$. Both uniform and pseudolikelihood proposals were used. Simulations reported in this Section, as well as Sections 5.3 and 5.4, were run on the high performance computing cluster Fox provided by the University of Oslo.
@@
Table 1: Results of simulations with sequentially arriving top-3 rankings. Values $\alpha$ and $d(\bm{\rho},\hat{\bm{\rho}})$ are Monte Carlo averages of the posterior means, with standard errors in parentheses.
Proposal Resampler $\alpha$ $d(\bm{\rho},\hat{\bm{\rho}})$ Time (minutes)
Pseudolikelihood Multinomial 0.292 (4e-04) 3.42 (0.023) 160.2 (0.8)
Residual 0.295 (4e-04) 3.23 (0.021) 167.1 (0.8)
Stratified 0.290 (4e-04) 3.68 (0.022) 165.9 (0.8)
Systematic 0.295 (4e-04) 3.70 (0.026) 158.1 (0.8)
Uniform Multinomial 0.294 (4e-04) 3.34 (0.021) 14.4 (0.1)
Residual 0.296 (4e-04) 3.33 (0.021) 15.2 (0.1)
Stratified 0.296 (5e-04) 3.40 (0.025) 15.8 (0.2)
Systematic 0.295 (4e-04) 3.61 (0.027) 17.2 (0.2)
@@
Table 1 shows simulation results after the final timepoint. The third column shows that the final estimate of $\alpha$ was close to the true value in all cases. With pseudolikelihood proposal, residual resampling gave the lowest posterior mean distance to the true ranking. With both proposals, systematic resampling gave the highest posterior mean distance to the true ranking. Overall, the performance of uniform and pseudolikelihood proposal seems comparable. Processing a single dataset on eight cores took on average between 14.4 and 17.2 minutes when using uniform proposal, and between 158.1 and 167.1 minutes with pseudolikelihood proposal. In contrast, obtaining 10,000 posterior draws using batch estimation with MCMC took on average five seconds.
@@
Figure 5: Number of particle filters per core and resampling probability for simulations with sequentially arriving top-3 rankings. Shaded regions are 95% confidence bands.
@@
Figure 5 shows the number of particle filters per core as a function of time (left) and the resampling probability per core as a function of time (right). The plots for pseudolikelihood proposal were almost identical, and are not shown. The curves were obtained by fitting generalized additive models (GAMs) (Wood,, 2017) with ten thin-plate regression splines (Wood,, 2003) as basis functions to the number of particle filters and a binary resampling indicator, respectively. A unit link function was used for the number of particles and a logit link for the resampling indicator, and smoothing was done with restricted maximum likelihood. The left plot shows that the number of particles grows close to linearly. This is expected from Theorem 1 of Andrieu et al., (2010), which implies that the number of particle filters must grow linearly with the number of observations for the acceptance rate to stay constant. The right plot shows that the resampling probability decreases with the number of timepoints. Since each resampling step is followed by one or more rejuvenation steps, this means that as rejuvenation becomes more computationally demanding due to more observations, it also becomes less frequent. This behavior agrees with what is expected from theory; in particular, Theorem 1 and Section 4.3 of Chopin, (2002) predicts that the time interval between each time when resampling is needed should increase geometrically in the total number of observations, which would produce an exponentially decaying curve (see also Proposition 17.1 in Chopin and Papaspiliopoulos,, 2020).
@@ 5.3 Pairwise Preferences
We next simulated consistent pairwise preference data by first generating complete rankings of $m=5$ items with $\bm{\rho}=(1,2,\dots,5)^{\prime}$ and $\alpha=0.3$ for 200 users, and then randomly selecting four implied pairwise preferences for each user. The users were assumed to arrive sequentially, one at each timepoint. The process was repeated 30 times, and each dataset was processed in parallel on 30 CPUs, each with $S=100$ particles. The initial number of particle filters was 20, residual resampling was used, and all other parameters were as described in the previous sections.
Across the 30 datasets, the average posterior mean of $\alpha$ at the final timepoint was $0.329$, with 95% Monte Carlo interval $(0.327,0.331)$. For comparison, the average posterior mean of $\alpha$ obtained by MCMC estimation with 20,000 post-burnin iterations on the same datasets was $0.308$ with 95% Monte Carlo interval $(0.306,0.310)$, suggesting that SMC2 was slightly positively biased in this case. The average wall time for computing the posterior for a simulated dataset was 2.0 minutes, compared to 7.3 seconds for batch estimation with MCMC.
Figure 6 (left) shows how the posterior expectation of $\alpha$ evolved as more data became available, with initial rapid fluctuations followed by a stabilization close to the true value of $0.3$ after about 100 timepoints. The center plot in Figure 6 shows the posterior probability that item 1 is preferred to item 2 in the modal ranking, and the right plot show the posterior probability that item 1 is preferred to item 3. As expected, the latter converges more quickly to one than the former. In particular, for three of the simulated datasets $P(\rho_{1}<\rho_{2})$ was below 0.5 even after the final timepoint, implying that items 1 and 2 were "flipped" in the posterior distribution, while $P(\rho_{1}<\rho_{3})$ was close to 1 for all datasets.